Math 042: Exam I Date: Feb 24, 2010

Name: Section:

1. Simplify: 7a^{3} − 2a^{2} − (5a^{2} − 6) Your answer: .

2. Solve for x: 5x − 9 = 2x − 16 Your answer: .

3. Solve for x: \frac{2}{3}x − 1 = \frac{3}{4}x Your answer: .

4. Simplify: \frac{(5a/7b^{2})}{(15/8ab)} Your answer: .

5. Evaluate: −a^{2} + 3b^{3} for a = 6, b = −3

(a) 45 (b) -45 (c) 117 (d) -117 6. Simplify: \frac{(3ab)^{2}}{9a^{2}b^{3}} Your answer: .

7. Solve the inequality: −2(y + 4) > 18 Your answer: .

8. Solve for c: f = \frac{9}{5}c + 32

(a) \frac{5(f−32)}{9} (a) \frac{5f−32}{9} (c) \frac{9}{5}f − 32 (d) \frac{5f}{9} − 32 9. Solve for x: cx + b = ax − d Your answer: .

10. 160% of what number is 80? Your answer: .

11. 72 is what percent of 90? Your answer: .

12. 6 times the sum of a number and its cube can be represented by

(a) 6(n + n^{3}) (b) 6(n + n)^{3} (c) 6 + n + n^{3} (d) 6n + n^{2} 13. The length of a rectangle is 5 less than twice its width. What can be used to represent its perimeter?

Your answer: .

14. An RCA employee earned $240 a week. If $32 is deducted for taxes, $25 for health benefits and $15 for pension, what percent of the salary is left?

(a) 30% (b) 70% (c) 72% (d) 168%

15. If $1100 is invested at t% and $700 is invested at (t − 2)% what can be used to represent the interest drawn in one year?

Your answer: .